### v1

# AdaptGCD: Multi-Expert Adapter Tuning for Generalized Category Discovery

Abstract
Different from the traditional semi-supervised learning paradigm that is constrained by the close-world assumption, Generalized Category Discovery (GCD) presumes that the unlabeled dataset contains new categories not appearing in the labeled set, and aims to not only classify old categories but also discover new categories in the unlabeled data. Existing studies on GCD typically devote to transferring the general knowledge from the self-supervised pretrained model to the target GCD task via some fine-tuning strategies, such as partial tuning and prompt learning. Nevertheless, these fine-tuning methods fail to make a sound balance between the generalization capacity of pretrained backbone and the adaptability to the GCD task. To fill this gap, in this paper, we propose a novel adapter-tuning-based method named AdaptGCD, which is the first work to introduce the adapter tuning into the GCD task and provides some key insights expected to enlighten future research. Furthermore, considering the discrepancy of supervision information between the old and new classes, a multi-expert adapter structure equipped with a route assignment constraint is elaborately devised, such that the data from old and new classes are separated into different expert groups. Extensive experiments are conducted on 7 widely-used datasets. The remarkable improvements in performance highlight the effectiveness of our proposals.
1 Introduction
While recent semi-supervised studies DBLP:conf/nips/SohnBCZZRCKL20 ; DBLP:conf/nips/BerthelotCGPOR19 ; DBLP:conf/nips/ZhangWHWWOS21 ; DBLP:conf/nips/LiWLYY0L23 ; DBLP:conf/nips/SunSL23 ; DBLP:conf/nips/Huang0Y0L23 show a substantial progression, they are typically constrained by the closed-world assumption, i.e., labels are provided for each category that the model needs to classify DBLP:conf/iclr/CaoBL22 . The constraint hinders the scalability of these methods. To relax this assumption, Generalized Category Discovery (GCD) DBLP:conf/cvpr/gcdVazeHVZ22 presumes that the unlabeled dataset contains new categories not appearing in the labeled set, as shown in Figure 1 (a). The goal of GCD is not only to classify old categories but also to discover new categories in the unlabeled data.
The primary challenge in GCD is learning the expressive representation applicable to both old and new classes DBLP:conf/cvpr/gcdVazeHVZ22 . Due to the scarcity of labels in GCD tasks, the model trained from scratch is prone to overfit the labeled data, consequently disrupting the generality of representation wang2024sptnet . As an alternative, the existing studies in GCD broadly finetune the self-supervised pretrained models, like DINO DBLP:conf/iccv/CaronTMJMBJ21 , to transfer the strong general knowledge from the pretrained backbone to the target GCD data. In this process, a key concern is to strike a balance between the generalization capacity from the upstream knowledge and the adaptability to the downstream GCD task, which is largely influenced by the fine-tuning strategy DBLP:journals/corr/abs-2403-14608 ; DBLP:conf/nips/LiuTMMHBR22 ; DBLP:conf/eccv/vptJiaTCCBHL22 ; DBLP:conf/acl/LiL20 .
Figure 1: (a) The description of the generalized category discovery (GCD) task. (b) The framework of our proposed AdaptGCD. It contains two critical modules: the multi-expert adapter (MEA) structure and the route assignment constraint. The MEA introduces multiple adapter experts and the route assignment constraint guides the allocation of these experts.
Up to now, two fine-tuning strategies are widely utilized in GCD: the partial tuning and the prompt tuning. The straightforward partial tuning strategy only updates the parameters from the last few blocks while the rest parameters in the backbone are frozen DBLP:conf/cvpr/gcdVazeHVZ22 . Such approaches potentially impair the generalization ability since the pretrained knowledge embodied in the backbone is disrupted. Conversely, prompt tuning methods DBLP:conf/eccv/vptJiaTCCBHL22 ; DBLP:conf/nips/WangC0H23 like PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 and SPTNet wang2024sptnet freeze all network parameters in the backbone and incorporate a learnable prompt at the input for training. Although effectively preserving the knowledge of pretrained network, unfortunately, it is demonstrated that the prompt-based approaches are not robust to bad data quality wang2024sptnet , and at the same time present limited adaptability when the downstream task differs from the upstream pretraining task DBLP:conf/eccv/vptJiaTCCBHL22 , thus making their performance on GCD task far from satisfactory.
Luckily, as one of the most characteristic fine-tuning approaches, the adapter tuning DBLP:conf/nips/adaptformerChenGTWSWL22 ; DBLP:conf/nips/RebuffiBV17 ; DBLP:conf/nips/LeiBBALZ0ZWLZC23 ; DBLP:conf/nips/Page-CacciaPSPR23 ; DBLP:conf/nips/ChengSXW0GS23 may provide a promising solution to the issues mentioned above. The adapter tuning leaves the pretrained model fixed and brings in a modest number of task-specific parameters via bottleneck adapter modules DBLP:journals/corr/gloraabs-2306-07967 . Obviously, on mechanism, the adapter naturally fits for the GCD task. It could preserve the pretrained knowledge nondestructively DBLP:conf/acl/WangTDWHJCJZ21 ; DBLP:journals/taslp/HouZWWQXS22 ; DBLP:conf/miccai/ZhangHZZZW23 , as well as provide flexibility for inserting domain knowledge of diverse downstream GCD tasks DBLP:conf/acl/Ye020 ; DBLP:conf/cvpr/LiLB22 ; DBLP:conf/aaai/XinDWLY24 . Surprisingly, the adapter has not yet attracted much attention on the GCD task. When introducing the adapter into GCD, an additional concern arises accordingly, i.e., the information imbalance between old and new classes within the adapter DBLP:conf/iclr/CaoBL22 . Due to the existence of supervised loss for labeled data, the adapter receives more supervision from old classes than the new ones, thereby resulting in a bias towards old classes. Under the adapter framework, how to render old classes and new classes non-interfering in the GCD task deserves further exploration.
In this study, we propose a novel multi-expert adapter tuning method AdaptGCD for generalized category discovery as shown in Figure 1 (b). Unlike the previous partial tuning or prompt-like methods, AdaptGCD integrates learnable adapters parallel to the feed-forward layer of the ViT block and keeps the parameters from backbone frozen to preserve the pretrained knowledge. Considering the discrepancy of supervision information between the old and new classes, the adapter is further extended to a Multi-Expert Adapter structure (MEA), which encompasses multiple adapter experts and employs a routing layer to schedule the usage of these experts. In order to guide the experts to specialize in data from old or new classes, a route assignment constraint is designed. In summary, the key contributions are listed as follows:
(i) Our proposal is the first work to introduce the adapter tuning into the GCD task. Such combination is elaborate, as it mechanistically allows preserving the general knowledge of pretrained backbone and meanwhile enhancing the adaptability to the downstream GCD task. We believe this could provide crucial insight for the future work of GCD.
(ii) To reduce the information interference between old and new classes, we employ a multi-expert adapter structure and design a route assignment constraint to separate the data from old and new classes into different expert groups.
(iii) We conduct extensive experiments on 7 popular datasets. Impressively, the adapter improves the GCD performance by a large margin, which highlights the applicability of adapter tuning on the GCD task. Furthermore, the results also confirm the effectiveness of multi-expert adapter structure with route assignment constraint.
2 Related Work
Generalized Category Discovery. Generalized Category Discovery (GCD, DBLP:conf/cvpr/gcdVazeHVZ22 ; DBLP:conf/nips/VazeVZ23 ; DBLP:conf/nips/RastegarDS23 ; DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 ; wang2024sptnet ; DBLP:conf/nips/BaiLWCMLZFWH23 ) aims to classify the unlabeled images across both old and new classes. The pioneering work DBLP:conf/cvpr/gcdVazeHVZ22 approaches the goal via refining the representation from the pretrained model with contrastive learning and assigning the classification results by semi-supervised $k$-means clustering. Some concurrent works DBLP:conf/cvpr/PuZS23 ; DBLP:conf/iccv/gmmgcdZhaoW023 ; DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 also follow the contrastive representation learning paradigm. Distinguished from these methods, the parametric approach SimGCD DBLP:conf/iccv/simgcdWenZQ23 introduces a learnable classifier and self-distillation to supervise the classifier, achieving boosting performance. Despite the excellent performance, most of these works DBLP:conf/cvpr/gcdVazeHVZ22 ; DBLP:conf/iccv/simgcdWenZQ23 adopt the simple fine-tuning strategy only updating the last few blocks of the vision transformer, which disrupts the pretrained knowledge. Hence, this work is devoted to seeking an effective fine-tuning strategy for the GCD task to preserve the priority and enhance adaptability to the downstream tasks simultaneously.
Fine-Tuning Strategy. The mainstream fine-tuning approaches includes partial tuning, prompt tuning and adapter tuning methods. Partial tuning methods freeze most of the backbone and finetune a small portion of parameters, like linear heads DBLP:conf/cvpr/IofinovaPKA22 , MLP heads DBLP:journals/corr/abs-2003-04297 , or several blocks of the backbone DBLP:conf/cvpr/HeCXLDG22 . Although this strategy is widely used in GCD, it still falls short of maintaining the network’s generalization capabilities. Prompt tuning DBLP:conf/eccv/vptJiaTCCBHL22 ; DBLP:conf/iccv/E2VPTHanWCCWQL23 prepends a set of learnable vectors to the input and only updates these prompts during fine-tuning. Despite their admirable performance, they present limited adaptability when a large gap exists between the downstream task and the upstream pretraining task DBLP:conf/eccv/vptJiaTCCBHL22 . Adapter tuning DBLP:conf/nips/adaptformerChenGTWSWL22 ; DBLP:journals/corr/abs-2302-08106 integrates shallow modules with the fixed backbone to protect the priority in pretrained models. Recent works DBLP:conf/acl/Ye020 ; DBLP:conf/cvpr/LiLB22 ; DBLP:conf/aaai/XinDWLY24 also show its flexibility to adapt to different downstream tasks. Considering the excellent mechanism of adapter tuning, this work further concentrates on the application and modification of the adapter-tuning-based strategies in GCD.
Mixture of Expert. The Mixture of Experts (MoE) is a machine learning technique employing route layer to segment a single task space into multiple subtasks DBLP:journals/corr/abs-2402-08562 ; DBLP:journals/corr/abs-2311-03285 ; DBLP:journals/corr/abs-2310-18339 ; DBLP:journals/corr/abs-2310-18547 ; DBLP:journals/corr/abs-2307-13269 . Each subtask is specifically managed by individual expert networks and the prediction result is aggregation of the solutions provided by these experts. In recent years, there appear some studies focusing on the fusion of the MoE with Adapters or LoRA DBLP:journals/corr/abs-2310-18339 ; dou2024loramoe . For instance, LoRAMoE dou2024loramoe introduces a MoE-style plugin and a localized balancing constraint to segregate world knowledge and downstream information into two distinct sets of experts. Inspired by LoRAMoE, the proposed approach also introduces multi-expert technology to handle data from new and old classes in diverse experts, thus reducing interference during the learning process.
3 Method
In this work, a novel AdaptGCD framework (as depicted in Figure 1) is proposed for tackling the generalized category discovery task. In the following, we would elaborate on our proposal.
3.1 Setting of Generalized Category Discovery
Assume that the training dataset is composed of the labeled and unlabeled parts defined as $\mathcal{D}^{l}=\{(\textbf{x}^{l}_{i},y^{l}_{i})\}\in\mathcal{X}\times\mathcal{Y}^{l}$ and $\mathcal{D}^{u}=\{(\textbf{x}^{u}_{i},y^{u}_{i})\}\in\mathcal{X}\times\mathcal{Y}^{u}$, where $\mathcal{Y}^{l},\mathcal{Y}^{u}$ are the label spaces of the labeled and unlabeled samples. In the GCD setting, $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$. Given $\mathcal{Y}^{n}=\mathcal{Y}^{u}\backslash\mathcal{Y}^{l}$, the categories in the $\mathcal{Y}^{l}$ are the “old” classes while those in $\mathcal{Y}^{n}$ are the “new” classes. The goal of GCD is to learn a model to categorize the samples from $\mathcal{D}^{u}$ in both the old and new classes.
3.2 Baselines: SimGCD
As a simple yet effective parametric approach, SimGCD DBLP:conf/iccv/simgcdWenZQ23 is adopted as the baseline of the proposed method. It incorporates two vital losses, including representation learning loss $\mathcal{L}_{\rm rep}$ and classification objective $\mathcal{L}_{\rm cls}$.
Parametric Classifier. SimGCD employs the parametric classifier in conjunction with the self-distillation paradigm for the learning of the classifier. Given the set of prototypes $\mathcal{C}=\{\textbf{c}_{1},...\textbf{c}_{K}\}$, where each one corresponds to a class and $K=|\mathcal{Y}_{u}|$ is the total number of categories. The prediction is determined by the cosine distance between the hidden features $\textbf{h}_{i}=f(\textbf{x}_{i})$ and the prototypes, that is ${p}_{i,k}=\frac{\exp\left(\frac{1}{\tau_{s}}(\textbf{h}_{i}/||\textbf{h}_{i}||_{2})^{\top}(\textbf{c}_{k}/||\textbf{c}_{k}||_{2})\right)}{\sum_{k^{\prime}}\exp\left(\frac{1}{\tau_{s}}(\textbf{h}_{i}/||\textbf{h}_{i}||_{2})^{\top}(\textbf{c}_{k^{\prime}}/||\textbf{c}_{k^{\prime}}||_{2})\right)}$, where $\tau_{s}$ indicates the temperature.
SimGCD Loss. For the representation loss, it integrates supervised contrastive learning DBLP:journals/corr/abs-2004-11362 for labeled samples with self-supervised contrastive learning DBLP:conf/icml/ChenK0H20 for all samples. Considering the classification loss, it is defined as the cross-entropy loss between the predictions and pseudo-labels or ground-truth labels. All in all, the objective for SimGCD $\mathcal{L}_{\rm gcd}$ is described as the sum of representation loss $\mathcal{L}_{rep}$ and classification loss $\mathcal{L}_{cls}$, i.e., $\mathcal{L}_{\rm gcd}=\mathcal{L}_{rep}+\mathcal{L}_{cls}$.
Partial Tuning. The effectiveness of the SimGCD framework largely stems from pretrained model. To take advantage of the generic representation from upstream tasks (like DINO DBLP:conf/iccv/CaronTMJMBJ21 ), the SimGCD framework only unfreezes the pretrained parameters of the last block for fine-tuning. However, as mentioned previously, this method might still result in information loss from the pretrained backbone, thereby reducing the generality of features.
3.3 Multi-Expert Adapter
In this part, the proposed approach integrates the adapter structure into the GCD framework, introducing a learnable additional architecture into the backbone while keeping the rest parameters in the backbone frozen during training. Then the basic adapter is further developed into the multi-expert adapter, allowing data under different supervisions to navigate through diverse paths.
Basic Adapter. The designation of the adapter in this work follows the basic architecture proposed by AdaptFormer DBLP:conf/nips/adaptformerChenGTWSWL22 , which replaces the Feed-Forward Network (FFN) in the Transformer with the AdaptMLP module. This module contains two parallel sub-branches, one is identical to the Feed-Forward block while the other is an learnable lightweight bottleneck module. The bottleneck module includes a down-projection layer with parameters $\textbf{W}_{\rm down}\in\mathbb{R}^{\hat{d}\times d}$, $\textbf{b}_{\rm down}\in\mathbb{R}^{1\times\hat{d}}$ and an up-projection layer $\textbf{W}_{\rm up}\in\mathbb{R}^{d\times\hat{d}}$, $\textbf{b}_{\rm up}\in\mathbb{R}^{1\times{d}}$, where $d$ and $\hat{d}$ denote the input dimension of FFN and middle dimension of bottleneck ($\hat{d}<d$). The down-projection layer and the up-projection layer are connected with the ReLU layer for non-linear properties and the entire bottleneck module is linked to the original FFN with the residual connection, as depicted in Figure 1. Specifically, for the features of the $i$-th sample fed into FFN of the $l$-th Transformer blocks $\widetilde{\textbf{x}}_{l,i}$, the forward process of bottleneck module to yield the adapted feature $\Delta{\textbf{x}}_{l,i}$ is formulated as follows:
$\Delta{\textbf{x}}_{l,i}=s\left({\rm ReLU}(\widetilde{\textbf{x}}_{l,i}\textbf{W}_{\rm down}^{\top}+\textbf{b}_{\rm down})\textbf{W}_{\rm up}^{\top}+\textbf{b}_{\rm up}\right),$ (1)
where $s$ is the scale factor. Then, the adapted features are fused with the original FFN branch by residual connection:
${\textbf{x}}_{l+1,i}=\overbrace{{\rm MLP}({\rm LN}(\widetilde{\textbf{x}}_{l,i}))+\widetilde{\textbf{x}}_{l,i}}^{\text{Original FFN}}+\overbrace{\Delta{\textbf{x}}_{l,i}}^{\text{Bottleneck}},$ (2)
where $\textbf{x}_{l+1,i}$ is the output of the $l$-th Transformer block, LN($\cdot$) is the layer normalization function. During fine-tuning, only the parameters of the bottleneck modules are updated and the rest are frozen.
Multiple-Expert Adapter. As previously mentioned DBLP:conf/iclr/CaoBL22 , data from older classes yields more supervisory information. It would result in the preference for the old class as the network receives more gradients from the old-class data than the new-class data. To counteract this implicit bias, the strategy of multi-expert technology is adopted to integrate with the adapter. This approach assigns different types of data to different expert adapters, therefore achieving separation for the data at the network level and alleviating the interference between new classes and old classes.
Multiple-expert adapter (MEA) consists of $T$ experts for handling different features and one route function to assign weights to these experts. More specifically, for the input feature $\widetilde{\textbf{x}}_{l,i}$, MEA leverages the following route function to estimate the weight $\boldsymbol{\omega}_{l,i}$ of multiple experts:
$\boldsymbol{\omega}_{l,i}=\left[\omega^{1}_{l,i},...\omega^{t}_{l,i},...,\omega^{T}_{l,i}\right]=$ ${\rm Softmax}(\widetilde{\textbf{x}}_{l,i}\textbf{W}_{\rm route}^{\top}/\tau_{r}),$ (3)
where $\textbf{W}_{\rm route}\in\mathbb{R}^{T\times d}$ denotes the trainable parameters in the route function and $\tau_{r}$ indicates the tempertures. The forward process of the bottleneck module is rectified as Eq.(4):
$\Delta{\textbf{x}}_{l,i}=s\sum_{t=1}^{T}\omega^{t}_{l,i}\left({\rm ReLU}(\widetilde{\textbf{x}}_{l,i}\textbf{W}_{\rm down}^{t}+\textbf{b}_{\rm down}^{t})\textbf{W}_{\rm up}^{t}+\textbf{b}_{\rm up}^{t}\right),$ (4)
where $\textbf{W}_{\rm down}^{t},\textbf{b}_{\rm down}^{t},\textbf{W}_{\rm up}^{t}$ and $\textbf{b}_{\rm up}^{t}$ denotes the parameters of the $t$-expert. By this, the multi-expert mechanism enables the network to develop powerful capabilities to flexibly handle the data from new and old classes. According to previous research DBLP:journals/corr/abs-2402-08562 , allocating more adapter experts in higher blocks enhances the effectiveness of models compared with inserting them in the lower blocks. Hence, the MEA structure is incorporated only in the last $P$ blocks out of the total $L$ blocks.
3.4 Route Assignment Constraint
In the multi-expert adapter, the route assignment constraint is required to supervise and control the route distribution. The assignment mainly focuses on two aspects: First, for all data, the load of all experts needs to be balanced to make full use of the resources of experts. Second, for data in old or new classes, the constraint assigns the corresponding experts to them so that the data can be well separated at the routing level. These two aspects correspond to the balanced load loss and the partial balanced load loss, which are introduced in this part.
Probability for Route Assignment. It is worthy noting that, in order to obtain the sample-wise route assignment probability, it is necessary to combine the gate vectors of different tokens for the same sample. Assume that $\boldsymbol{\omega}_{l,i,v}$ denotes the gate probability from the $v$-th token of the $i$-th sample, within the $l$-th blocks, then the route assignment probability for the $i$-th sample ${\boldsymbol{\omega}}_{l,i}$ can be formulated as ${\boldsymbol{\omega}}_{l,i}={\rm Softmax}\left(\frac{1}{V}\sum_{v=1}^{V}\boldsymbol{\omega}_{l,i,v}/\tau_{g}\right)$, where $V$ denotes the number of tokens and $\tau_{g}$ is the temperature, and ${\rm Softmax}(\cdot)$ is utilized to sharpen the gate vectors.
Balanced Load Loss. The balanced load loss is designed to ensure the maximal usage of diverse experts. Previous research on mixture-of-expert DBLP:journals/corr/loramoeabs-2312-09979 demonstrated a tendency towards a “winner-takes-all” DBLP:journals/corr/loramoeabs-2312-09979 situation, which indicates that few experts are frequently activated and domain the learning process, resulting in the waste of some experts. To mitigate this, in this study, the mean distribution of experts in the $l$-th layer ${\boldsymbol{\omega}}_{l}$, averaging the route assignment probabilities across all samples, is aligned to a uniform distribution $I$, where $I(t)=1/T,t=1,...T$. The formulation is depicted in Eq. (5):
${\boldsymbol{\omega}}_{l}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}{\boldsymbol{\omega}}_{l,i},\quad\mathcal{L}_{bl}=\sum_{l=L-P+1}^{L}\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}_{l}||{I}\right),$ (5)
where $\mathcal{B}$ indicates the mini-batch and $\mathbb{D}_{\rm KL}(\cdot||\cdot)$ indicates the Kullback-Leibler divergence.
Partial Balanced Load Loss. Considering the imbalance of supervision between new classes and old classes, we propose a partial balanced load loss to separate the new-class and old-class data into different experts and reduce their interference. Since the exact class label of the samples is not determined during training, the pseudo labels are utilized as a substitute. As shown in Eq. (6), the pseudo-labels of the samples are obtained according to the predictions.
[…]
Then the route assignment probabilities of the experts for the old and new classes are estimated as ${\boldsymbol{\omega}}^{old}_{l}$ and ${\boldsymbol{\omega}}^{new}_{l}$ in Eq. (7).
${\boldsymbol{\omega}}^{old}_{l}=\frac{\sum_{i}{\boldsymbol{\omega}}_{l,i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{l})}{\sum_{i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{l})},\quad{\boldsymbol{\omega}}^{new}_{l}=\frac{\sum_{i}{\boldsymbol{\omega}}_{l,i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{n})}{\sum_{i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{n})}.$ (7)
We manually specify the expert groups for the old and new classes beforehand and denote the expert groups as $\mathcal{T}^{old}$ and $\mathcal{T}^{new}$, respectively. For instance, the first four experts are assigned to $\mathcal{T}^{old}$ and the remaining experts are naturally divided into $\mathcal{T}^{new}$. Then the target route distribution probability for the old classes $I^{old}\in\mathbb{R}^{T}$ and new classes $I^{new}\in\mathbb{R}^{T}$ is established as follows:
$I^{old}(t)=\begin{cases}\frac{1}{|\mathcal{T}^{old}|}&t\in\mathcal{T}^{old}\\ 0&t\in\mathcal{T}^{new}\\ \end{cases},I^{new}(t)=\begin{cases}0&t\in\mathcal{T}^{old}\\ \frac{1}{|\mathcal{T}^{new}|}&t\in\mathcal{T}^{new}\\ \end{cases}.$ (8)
These two distribution functions respectively describe the target assignment distribution of data in the old and new data. Then, as shown in Eq. (9), the Kullback-Leibler divergence is adopted to align ${\boldsymbol{\omega}}^{old}_{l}$ and ${\boldsymbol{\omega}}^{new}_{l}$ with the predefined target $I^{old}$ and $I^{new}$:
$\mathcal{L}_{pbl}=\sum_{l=L-1+P}^{L}\left(\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}^{old}_{l}||{I}^{old}\right)+\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}^{new}_{l}||{I}^{new}\right)\right).$ (9)
In the final step, the route assignment loss $\mathcal{L}_{ra}$ for the AdaptGCD is collected as the weighted sum of the two losses, i.e. $\mathcal{L}_{ra}=\beta\mathcal{L}_{bc}+\alpha\mathcal{L}_{pbl}$, where $\alpha,\beta$ are the balancing factors.
Overall Loss. The overall loss of the AdaptGCD is collected as the sum of SimGCD loss and the route assignment loss, as formulated in Eq. (10):
$\mathcal{L}_{overall}=\mathcal{L}_{gcd}+\mathcal{L}_{ra}.$ (10)
4 Experiments
4.1 Experimental Setup
Datasets. The effectiveness of the proposed method is validated on three characteristic benchmark: the generic image recognition benchmark including CIFAR10/100 krizhevsky2009learning and ImageNet-100 DBLP:conf/eccv/TianKI20 , the semantic shift benchmark including CUB-200 welinder2010caltech , Stanford Cars krause20133d , and FGVC-Aircraft DBLP:journals/corr/MajiRKBV13 and the harder long-tail dataset Herbarium 19 tan2019herbarium . Following SimGCD DBLP:conf/iccv/simgcdWenZQ23 , a subset of all the classes $\mathcal{Y}^{l}$ is sampled as the old classes while 50% of the images from the old classes are selected to construct the labeled set $\mathcal{D}^{l}$. The rest of the images are collected as the unlabeled set $\mathcal{D}^{u}$.
Implementation details and evaluations. Following GCD DBLP:conf/cvpr/gcdVazeHVZ22 , the proposed method is trained with a ViT-B/16 backbone pre-trained with DINO. The output of [CLS] tokens is utilized as the feature of the image. For the training details, the AdaptGCD is trained with a batch size of 128 for 200 epochs with an initial learning rate of 0.1 decayed with a cosine schedule on each dataset. For the adapter configuration, we utilize 8 experts with the bottleneck dimension $\hat{d}$ set as 64. The number of adapted blocks $P$ is assigned to be 6 on CUB-200, Scars and CIFAR10. As for the Aircraft, CIFAR100, Herbarium19 and Imagenet100, which present greater challenges for the classification, $P$ is designated to be 8. For expert assignments, 4 experts are utilized as old-class experts while another 4 experts are selected as new-class experts, i.e. $|\mathcal{T}^{old}|=|\mathcal{T}^{new}|=4$. The performance is evaluated with the clustering accuracy for old, new and all classes as DBLP:conf/cvpr/gcdVazeHVZ22 . It is also noted that we conduct the experiments in Table 1 and 2 for 3 independent runs and report the average.
Table 1: Evaluation results on the generic image recognition benchmark. Bold values represent the best results, while underlined values denote the second best results.
Methods CIFAR10 CIFAR100 ImageNet100
All Old New All Old New All Old New
$k\mbox{-}means$ DBLP:conf/soda/ArthurV07 83.6 85.7 82.5 52.0 52.2 50.8 72.7 75.5 71.3
RankStats+ DBLP:journals/pami/HanREVZ22 46.8 19.2 60.5 58.2 77.6 19.3 37.1 61.6 24.8
UNO+ DBLP:conf/iccv/FiniSLZN021 68.6 98.3 53.8 69.5 80.6 47.2 70.3 95.0 57.9
ORCA DBLP:conf/iclr/CaoBL22 81.8 86.2 79.6 69.0 77.4 52.0 73.5 92.6 63.9
GCD DBLP:conf/cvpr/gcdVazeHVZ22 91.5 97.9 88.2 73.0 76.2 66.5 74.1 89.8 66.3
DCCL DBLP:conf/cvpr/PuZS23 96.3 96.5 96.9 75.3 76.8 70.2 80.5 90.5 76.2
SimGCD DBLP:conf/iccv/simgcdWenZQ23 97.1 95.1 98.1 80.1 81.2 77.8 83.0 93.1 77.9
PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 97.9 96.6 98.5 81.2 84.2 75.3 83.1 92.7 78.3
SPTNet wang2024sptnet 97.3 95.0 98.6 81.3 84.3 75.6 85.4 93.2 81.4
AdaptGCD 97.9 95.5 99.0 84.0 86.2 79.7 86.4 94.9 82.1
Table 2: Evaluated results on the semantic shift benchmark and Herbarium19. Bold values represent the best results, while underlined values represent the second best results.
Methods CUB-200 Scars Aircraft Herbarium19
All Old New All Old New All Old New All Old New
$k\mbox{-}means$ DBLP:conf/soda/ArthurV07 34.3 38.9 32.1 12.8 10.6 13.8 16.0 14.4 16.8 13.0 12.2 13.4
RankStats+ DBLP:journals/pami/HanREVZ22 33.3 51.6 24.2 28.3 61.8 12.1 26.9 36.4 22.2 27.9 55.8 12.8
UNO+ DBLP:conf/iccv/FiniSLZN021 35.1 49.0 28.1 35.5 70.5 18.6 40.3 56.4 32.2 28.3 53.7 14.7
ORCA DBLP:conf/iclr/CaoBL22 35.3 45.6 30.2 23.5 50.1 10.7 22.0 31.8 17.1 35.4 51.0 27.0
GCD DBLP:conf/cvpr/gcdVazeHVZ22 51.3 56.6 48.7 39.0 57.6 29.9 45.0 41.1 46.9 20.9 30.9 15.5
DCCL DBLP:conf/cvpr/PuZS23 63.5 60.8 64.9 43.1 55.7 36.2 - - - - - -
SimGCD DBLP:conf/iccv/simgcdWenZQ23 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8 44.0 58.0 36.4
PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 62.9 64.4 62.1 50.2 70.1 40.6 52.2 52.2 52.3 37.0 52.0 28.9
SPTNet wang2024sptnet 65.8 68.8 65.1 59.0 79.2 49.3 59.3 61.8 58.1 43.4 58.7 35.2
AdaptGCD 68.8 74.5 65.9 62.7 80.6 54.0 57.9 65.2 54.3 45.7 60.6 37.7
4.2 Comparison with State-of-the-Arts
Evaluation on generic image recognition datasets. AdaptGCD is compared with the previous advanced methods and some concurrent works (i.e., DCCL DBLP:conf/cvpr/PuZS23 , PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 , SimGCD DBLP:conf/iccv/simgcdWenZQ23 and SPTNet wang2024sptnet ). The results are shown in Table 1. It is observed that our AdaptGCD manifests superior performance across three general benchmarks. Compared to the baseline SimGCD, AdaptGCD achieves improvements of 0.8%, 3.9%, 3.1% for “All” classes on three datasets. When juxtaposed with other concurrent methods, our method advances on all the metrics except for the “Old” classes on CIFAR10. On CIFAR10 datasets, AdaptGCD prioritizes a balance between the old and new categories so that it outperforms the GCD by 10.8% in “New” classes, at a cost of 2% dip in the “Old” classes. It is also noted that the most advanced prompt-based method, SPTNet, underperforms its baseline SimGCD on several metrics on CIFAR10 and CIFAR100 datasets since the low-resolution images hinder the utility of spatial prompts. The promising performance achieved by our AdaptGCD further validates that the adapter-like methods could exhibit excellent adaptability to data with varying qualities than the prompt-like rivals.
Evaluation on fine-grained datasets. Table 2 presents the results on fine-grained datasets including the semantic shift benchmark and the long-tail dataset Herbarium19. Due to the large intra-class and small inter-class variations for data, it is more difficult to discover fine-grained categories than the generic classes. Despite the difficulty, AdaptGCD shows consistently superior performance on the CUB-200, Scars and Herbarium19 datasets, achieving an average improvement of 3% on “All” classes in comparison with the state-of-the-art results. Even within the Aircraft dataset, it achieves the second best scores with only 1.4% lag on “All” classes. These observations indicate the efficacy of the proposed method across fine-grained datasets.
4.3 Ablation Study
To investigate the impact of different modules, we conduct the the extensive ablation studies on the semantic shift benchmark following wang2024sptnet ; DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 . The results are reported in Table 3.
Table 3: Ablation study on the semantic shift benchmark
[…]
All Old New All Old New All Old New
(1) SimGCD - - $\mathcal{L}_{gcd}$ 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8
(2) SimGCD + MEA 1 64 $\mathcal{L}_{gcd}$ 67.5 75.3 63.5 59.8 77.4 51.2 55.2 65.0 50.3
(3) SimGCD + MEA 1 512 $\mathcal{L}_{gcd}$ 68.0 76.1 63.9 61.3 77.5 53.5 57.7 66.9 53.1
(4) SimGCD + MEA 8 64 $\mathcal{L}_{gcd}$ 68.5 74.9 65.3 61.4 79.4 52.7 57.3 65.3 53.2
(5) SimGCD + MEA 8 64 $\mathcal{L}_{gcd}+\mathcal{L}_{bl}$ 68.6 75.0 65.4 62.0 80.9 52.8 57.7 65.3 53.9
(6) SimGCD + MEA 8 64 $\mathcal{L}_{gcd}+\mathcal{L}_{bl}+\mathcal{L}_{pbl}$ 68.8 74.5 65.9 62.7 80.6 54.0 57.9 65.2 54.3
Influence of adding adapter structure. Firstly, we have validated the role of a single adapter as illustrated in (2). Given that the bottleneck dimension is set as 64, the count of learnable parameters in the backbone is less than 0.8M, which is significantly less than the number of parameters in the last block (approximately 7M). Despite the less trainable parameters in the backbone, involving an adapter still achieves remarkable accuracy boost of 7.2%, 6.0% and 1.0% for “All" classes on the CUB-200, Scars and Aircraft datasets. These findings affirm the adapter’s efficacy for GCD task.
Influence of multi-expert adapter and route assignment constraint. To validate the effectiveness of the multi-expert adapter (MEA), the method implemented with MEA (in (4)) is compared with two alternate single adapter approaches. One incorporates an equivalent bottleneck dimension, as shown in (2), while the other uses the setup with a comparable parameter count, denoted in (3). With an identical bottleneck dimension, the results reveal that the MEA outperforms the single expert method, thereby establishing the MEA’s superiority. On the other hand, when the number of learnable parameters is similar (see (3) vs. (4)), the MEA does not show obvious advantages. Nevertheless, with the assistance of route assignment constraints (see (5),(6)), the methods incorporating MEA surpass the single adapter method (3) in terms of “All" classes across three datasets. This statement emphasizes the significant role that the route assignment constraint holds in the MEA training process. It is worth noting that the improvements derived from the application of MEA are not only due to the increased number of learnable parameters but also contribute to the separation of different data via route allocations, enabled by the constraint. It’s also observed that the $\mathcal{L}_{pbl}$ enhances the performance of “New” classes at the expense of a slight reduction in “Old” classes (see (6) vs. (5)). This observation also validates our motivation to alleviate the bias towards old classes.
4.4 Discussion
Figure 2: Sensitivity analysis on three critical hyper-parameters. (a) bottleneck dimension $\hat{d}$, (b) the number of adapted blocks $P$, (c) expert count $T$.
Sensitivity of three key hyper-parameters of the adapter. We evaluated the impact of three key hyper-parameters of the adapter: bottleneck dimension $\hat{d}$, number of adapted blocks $P$ and expert count $T$, and show the results in Figure 2. First, we conduct experiments with models employing single adapters across different $\hat{d}$ and report the results in Figure 2 (a). Remarkably, when the bottleneck dimension is relatively small, there is a noticeable rise in performance with the increase in the bottleneck dimension. As the dimension further elevates, performance fluctuates within a certain range due to the overfitting caused by the excessive amount of parameters. A similar trend is also observed as the increase in the number of adapted blocks $P$ in Figrue 2 (b), i.e., as $P$ is large enough, inserting more adapters in the bottle blocks yields additional resources but limited performance gain. It also validates the effectiveness of only inserting adapters in parts of blocks. For the high performance of a single adapter, the impact of $T$ in Figure 2 (c) is not so noticeable, but the increment in expert count still results in a performance boost for “All” classes, further corroborating the effectiveness of multi-expert adapters.
Figure 3: Attention visualization for 12 heads in the last blocks of backbone on CUB-200. “DINO” represents the results from pretrained backbone without additional training while “SimGCD” and “AdaptGCD” are the those from models trained via SimGCD and AdaptGCD.
Visualization of the attended maps. To better appreciate the advantages of the AdaptGCD, we extracted the attention maps corresponding to the [CLS] token from different heads in the last block of the ViT, highlighting the top 10 most notable patches in red for each head in Figure 3. Upon observation, it is evident that compared to SimGCD and DINO, AdaptGCD focuses more on the foreground and exhibits more varied attention regions in different heads (in heads 3/7/11/12), showing the superior adaptability of AdaptGCD to the CUB-200 datasets. It is also worth noting that in regions like heads 3/7/11, SimGCD fails to inherit the correct attention on the foreground from the DINO model while AdaptGCD refines the attention regions based on the DINO model. These findings further demonstrate the motivation to protect the pretrained information for generalization.
Figure 4: (a) Visualization of the route assignment probability for samples predicted as old and new classes in 8 experts. (b) Confusion maps for the old and new classes.
Visualization of route assignment probability. To verify the role of the route assignment constraint, we also visualize the assignment probabilities for samples predicted as old and new classes under the original SimGCD loss, the loss with $\mathcal{L}_{bl}$, and the loss with $\mathcal{L}_{bl}+\mathcal{L}_{pbl}$ on Aircraft dataset in Figure 4 (a). Without any additional loss, the “winner-takes-all” phenomenon occurs and the imbalance is greatly alleviated via $\mathcal{L}_{bl}$. Note that a significant portion of samples predicted as old or new classes are distributed to their corresponding experts with the addition of $\mathcal{L}_{bl}$ and $\mathcal{L}_{pbl}$. It validates the effectiveness of $\mathcal{L}_{pbl}$ to separate the data from old and new classes on the route level.
Confusion matrix of the model trained with balanced loss. In addition, to further investigate the role of the route assignment constraint in mitigating the bias towards old classes, we also visualized the confusion matrix of the corresponding method in Figure 4 (b). It is observed that with the constraint, the amount of new-class data misclassified as the old classes decreases, thus substantiating the effectiveness of the constraint in alleviating the preference for the old classes.
5 Conclusion
In this study, we propose a novel AdaptGCD in generalized category discovery, which is the pioneering work to merge adapter tuning into the GCD framework. It integrates learnable adapters with the fixed backbone to adapt to the downstream GCD task as well as preserve the pretrained priority. Additionally, considering the imbalance between the supervision of the old and new classes, we present a multi-expert adapter structure and a route assignment constraint to separate the data from old and new classes into different expert groups, further reducing mutual interference. The efficacy of the proposed strategy is demonstrated in 7 datasets, including 3 generic datasets 3 fine-grained datasets, and a long-tail dataset. Looking forward to future work, a promising direction is to explore more effective fine-tuning strategies in the GCD tasks, including the LoRA DBLP:conf/iclr/HuSWALWWC22 , RepAdapter DBLP:journals/corr/abs-2302-08106 , or the combination of different strategies like GLoRA DBLP:journals/corr/gloraabs-2306-07967 . In terms of limitations, the proposed method introduces additional parameters, which consume extra computational cost during training and inference.